# Multimodal Graph Benchmark

### v1

@@ 4 Evaluation Framework
To enable a comprehensive evaluation on multimodal graph data, MM-Graph standardizes the GNN architectures, feature encoders, and dataloaders and evaluators used across all datasets.
@@ 4.1 Graph Neural Networks
Conventional GNNs. We standardizes GCN, SAGE, MLP for both link prediction and node classification. Since BUDDY is specifically designed for link prediction, so we do not report its performance on node classification.
Multimodal GNNs. While GNNs have shown strong performance on tasks involving graphs with unimodal features, there has been limited work on developing GNNs that can effectively handle multimodal graph data. In order to evaluate the effectiveness of GNNs on multimodal graphs, we adapt two recent architectures from the recommendation systems domain to our standardized multimodal graph benchmarks:
MMGCN is a GCN-based framework that leverages information interchange between users and micro-videos in multiple modalities to refine their modality-specific representations and model users’ fine-grained preferences. It constructs separate user-item graphs for each modality. The information interchange of users and items in each modality is encoded using GNNs. And modality representations are fused in the prediction layer to predict possible recommendations [31]. For LP, we use the most common dot product decoder to decode possible links. For NC, we stack a 3-layer MLP as transform representations to a number of dimensions equal to the number of node classes.
MGAT extends the popular GAT architecture to multimodal settings. It learns modality-specific node representations which are then combined using a cross-modal attention layer  [28]. This allows MGAT to weigh different modalities based on their importance. Similar to MMGCN, we decode links using dot product for LP and perform NC through a 3-layer MLP.
@@ 4.2 Feature Encoders
Text encoders. For text encoders, we select CLIP  [24], T5  [25] and ImageBind  [8]. T5 is selected as the STOA text embedding models, while CLIP and ImageBind are selected so that the output text representations aligns with visual representations.
Visual encoders. For visual encoders, we select CLIP  [24], ViT  [4], ImageBind  [8] and Dinov2  [22]. ViT and Dinov2 are selected as two different types of visual encoders, while ViT are explicitly trained with supervision, Dinov2 learns robust visual features without supervision. CLIP and ImageBind are selected so that the output text representations aligns with visual representations. Specifically, Imagebind has potentials to extend multimodal graph learning to other modalities, such as audio and video thanks to its ability to bind all modalities into one embedding space. Studies on various feature encoders give insights on design choices of multimodal graph learning, such as (1) how important it is to align multiple modalities into a unified embedding space, (2) if supervised feature encoders work better than unsupervised feature encoders. Note that feature encoders are only used to extract features for GNNs and are never trained upon.
Table 3: Link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP. *OOM = out of GPU memory. Highlighted box indicates the best performing combination for each dataset.
@@
MMGCN CLIP 31.96 $\pm$0.10 16.35 $\pm$0.11 68.46 $\pm$0.08 22.20 $\pm$0.05 10.76 $\pm$0.1 46.62 $\pm$0.12 OOM OOM OOM
ViT T5 30.33 $\pm$0.03 15.01 $\pm$0.05 66.41 $\pm$0.11 19.45 $\pm$0.34 9.22 $\pm$0.20 40.49 $\pm$0.61 OOM OOM OOM
ImageBind 31.74 $\pm$0.21 16.45 $\pm$0.13 67.39 $\pm$0.74 24.72 $\pm$0.19 12.47 $\pm$0.09 51.32 $\pm$0.56 OOM OOM OOM
Dinov2 T5 30.04 $\pm$0.27 14.98 $\pm$0.07 64.56 $\pm$0.56 21.77 $\pm$0.23 10.47 $\pm$0.12 45.81 $\pm$0.52 OOM OOM OOM
MGAT CLIP 27.56 $\pm$0.30 13.55 $\pm$0.29 60.21 $\pm$0.21 21.38 $\pm$0.23 10.39 $\pm$0.22 44.60 $\pm$0.36 OOM OOM OOM
ViT T5 30.15 $\pm$0.34 15.28 $\pm$0.34 64.84 $\pm$0.41 20.59 $\pm$0.41 9.79 $\pm$0.30 43.44 $\pm$0.76 OOM OOM OOM
ImageBind 30.15 $\pm$0.12 15.50 $\pm$0.05 64.20 $\pm$0.43 22.13 $\pm$0.27 10.96 $\pm$0.15 45.84 $\pm$0.57 OOM OOM OOM
Dinov2 T5 28.91 $\pm$0.09 14.47 $\pm$0.18 62.11 $\pm$0.22 21.42 $\pm$0.13 10.38 $\pm$0.13 44.11 $\pm$0.50 OOM OOM OOM
@@ 4.3 Dataloader and Evaluators
MM-Graph provides a standardized dataloader and evaluators implementation for all five datasets. The dataloader is built upon PyTorch’s DataLoader class and the Deep Graph Library (DGL) . It efficiently handles the creation and transmission of graph samples, allowing for mini-batch training of GNNs. The evaluator is based on OGB’s evaluator and performs standardized and reliable evaluation.
The dataloader supports various sampling strategies, such as neighbor sampling and layer-wise sampling, to scale the training process to large graphs. It also enables distributed training using PyTorch’s distributed data parallel (DDP) for faster training on multiple GPUs.
By standardizing the GNNs, feature encoders, dataloader and evaluators, MM-Graph ensures a fair and comprehensive evaluation of graph learning algorithms on multimodal graph data.
@@ 5 Empirical Analysis
We conduct a comprehensive evaluation of our proposed benchmark on both link prediction and node classification tasks. Our experiments aim to answer the following research questions:
(i) Are multimodal-GNNs better than conventional GNNs on MM-Graph?
(ii) What is the best approach for encoding features? Is alignment between modalities necessary?
(iii) How much performance gain can be achieved by using multimodal features compared to text features alone?
To ensure fair and rigorous comparisons between various feature encoders and GNNs, we adopt consistent experimental settings across all experiments. We perform automatic hyperparameter tuning using Optuna [1] to optimize the performance of each model. Detailed experimental setup and hyperparameters are provided in the Appendix C.
Figure 2: Feature alignment is important. We compare the performance of various feature encoders, find aligned features, e.g. CLIP and Imagebind result in much better performance compared with unaligned features on Amazon-Sports and Amazon-Cloth. Among them, Imagebind performs the best across backbones. This indicates the importance of using aligned features on these datasets.
@@ 5.1 Results
The detailed link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP is shown in Table  3. The detailed node classification results on is shown in Table  4. Here are our findings:
Multimodal GNNs perform poorly on the proposed MM-Graph datasets. As shown in Fig  5.1, we observe that MMGCN and MGAT, the two multimodal GNNs specifically designed to process multimodal input data does not perform better than conventional GNN models such as SAGE. As MMGCN and MGAT mostly just perform message passing and aggregation on each modality separately and then fuse the information across modality together at last, we hypothesize that it is likely to do the insufficient fusion of modalities, and calls for better model designs of multimodal structured data.
Feature Encoders (Alignment). CLIP and ImageBind are two feature encoders which can map features of various modalities to the same embedding space while T5ViT and T5Dinov2 uses SOTA embedding methods for each modality separately, without any specific alignment layers. Interesting, we found the results to vary across datasets. As shown in Figure  2, we found that for Amazon-Sports and Amazon-Cloth, the two amazon datasets, aligned multimodal features performs consistently better than unaligned features. Specifically, we also found that ImageBind consistently gives the best results across backbones. On Goodreads-LP, we found that T5Dinov2 performs best most of the times, this indicates that Dinov2 excels other vision models in the Optical Character Recoginition (OCR) task.
Table 4: Node classification results on Goodreads-NC and Ele-Fashion. *OOM = out of GPU memory.
@@
MMGCN CLIP CLIP 86.10 $\pm$0.50 OOM
T5 ViT 82.39 $\pm$0.30 OOM
ImageBind ImageBind 86.21 $\pm$0.94 OOM
T5 Dinov2 85.53 $\pm$0.33 OOM
MGAT CLIP CLIP 84.66 $\pm$0.29 OOM
T5 ViT 84.01 $\pm$0.08 OOM
ImageBind ImageBind 86.12 $\pm$0.08 OOM
T5 Dinov2 84.54 $\pm$0.27 OOM
@@
Figure 3: Multimodal GNNs underperforms conventional GNNs. We compare the best performance of multimodal GNNs (MMGCN/MGAT) and conventional GNNs (SAGE, GCN, BUDDY). Conventional GNNs consistently perform better across datasets, which justifies the importance of building MM-Graph  and calls for better multimodal GNN designs.
[…]
5.2 Challenges and Future Directions.
First, existing multimodal GNNs are all based on full-batch training schemes and it would go out of GPU memory when the graph is large, e.g. Goodreads-NC and Goodreads-LP. Mini-batch multimodal GNNs are worth exploring in the future. Second, we find that it is hard to find a graph dataset with audio/video features with open-source access, even though music/video recommendation is an task of great importance in real-world deployed systems. We believe the study is similar to what is conducted here and the ImageBind features we use can be easily extended to these scenarios. However, integrating multimodal graphs with audio/video features still require future exploration.
Figure 4: Multimodal features are helpful for graph learning. Multimodal features, performs better than text features across datasets and tasks, which justifies the necessity of introducing multimodal graph datasets.
@@ 6 Conclusion
In summary, this paper introduces the Multimodal Graph Benchmark (MM-Graph), a comprehensive benchmark for multimodal graph data. MM-Graph consists of five diverse datasets spanning various domains and tasks, featuring rich textual and visual node features. To enable a fair and extensive evaluation, MM-Graph standardizes the GNN architectures, feature encoders, dataloaders and evaluators used across all datasets. The GNNs include both unimodal and multimodal architectures, while the feature encoders leverage state-of-the-art models for capturing semantic information from text and images. Experiments on MM-Graph datasets provide valuable insights into the performance of different GNN and feature encoder combinations in multimodal graph learning settings. By offering a unified evaluation framework, MM-Graph aims to facilitate research on multimodal graph representation learning and drive the development of more powerful and robust graph learning algorithms for real-world applications.
Social impacts. MM-Graph has the potential to significantly advance research on multimodal graph representation learning, which can enable novel applications that positively transform various aspects of society. Many real-world domains, such as healthcare, social networks, and urban computing, involve multimodal graph data. By providing a comprehensive evaluation framework, MM-Graph can accelerate the development of more powerful graph learning algorithms for these domains. MM-Graph also promotes reproducible research by offering standardized benchmarks, fostering collaboration within the research community. Ultimately, MM-Graph aims to unlock the value of multimodal graph data to enable impactful real-world applications. However, as with all technologies of this nature, it is important to consider how it can be used responsibly and with responsible development in mind. There are risks of these algorithms perpetuating or amplifying societal biases if the underlying graph data contains biased patterns. Additionally, misuse of multimodal graph can cause disinformation and targeted manipulation. It is therefore crucial that researchers using MM-GRAPH prioritize the responsible development of multimodal graph algorithms in line with ethical principles.

### v2

@@ 4 Evaluation Framework
To enable a comprehensive evaluation on multimodal graph data, MM-Graph standardizes the GNN architectures, KGEs, feature encoders, dataloaders and evaluators used.
@@ 4.1 Graph Neural Networks
Conventional GNNs. We standardize GCN, SAGE, MLP for both link prediction and node classification. Since BUDDY is specifically designed for link prediction, we do not report its performance on node classification. For all unimodal GNNs, text and image embeddings are first concatenated together and then passed through GNNs.
Multimodal GNNs. While GNNs have shown strong performance on tasks involving graphs with unimodal features, there has been limited work on developing GNNs that can effectively handle multimodal graph data. To evaluate the effectiveness of GNNs on multimodal graphs, we adapt two recent architectures from the recommendation systems domain to our standardized multimodal graph benchmarks:
(1) MMGCN constructs separate user-item graphs for each modality. The information interchange of users and items in each modality is encoded using GNNs. Modality representations are fused in the prediction layer to predict possible recommendations [43]. For link prediction (LP), we use the most common dot product decoder to decode possible links. For node classification (NC), we stack a 3-layer MLP to transform representations to a number of dimensions equal to the number of node classes.
(2) MGAT extends the popular GAT architecture and learns modality-specific node representations which are then combined using a cross-modal attention layer [39]. This allows MGAT to weigh different modalities based on their importance. Similar to MMGCN, we decode links using dot product for LP and perform NC through a 3-layer MLP.
4.2 Knowledge Graph Embeddings
MM-Graph includes two state-of-the-art multimodal Knowledge Graph Embeddings (KGEs): MoSE [49] and VISTA [20]. MoSE introduces modality-split relation embeddings for each modality, moving away from a single modality-shared embedding, which effectively mitigates modality interference. VISTA innovatively incorporates visual and textual representations of entities and relations through sophisticated transformers, utilizing entity encoding, relation encoding, and triplet decoding. Since VISTA’s architecture scores entities rather than triples, we report performance using the standard 1vsAll setting instead of hard negatives [20, 34].
Note that we do not report Graph Neural Network (GNN) performance for Knowledge Graph Completion (KGC), as prior research has demonstrated that GNN architectures can inadvertently conflate scoring functions in Knowledge Graph Embedding, potentially degrading overall performance [21].
@@ 4.3 Feature Encoders
Text encoders. For text encoders, we select CLIP [31], T5 [32], and ImageBind [10]. T5 is chosen as the state-of-the-art (SOTA) text embedding model, while CLIP and ImageBind are selected to ensure that the output text representations align with visual representations. Specifically, CLIP provides a multi-modal embedding approach that jointly trains text and image encoders to maximize cosine similarity between text and image embeddings. ImageBind extends this concept by enabling cross-modal representation learning across multiple modalities.
Table 5: Link prediction results on Amazon-Sports, Amazon-Cloth and Goodreads-LP. Highlighted box indicates the best performing combination for each dataset. Conventional GNNs such as SAGE performs best across datasets. Aligned feature embeddings, e.g. CLIP and ImageBind outperforms unaligned features.
@@
MMGCN CLIP 31.96 $\pm$0.10 16.35 $\pm$0.11 68.46 $\pm$0.08 22.20 $\pm$0.05 10.76 $\pm$0.1 46.62 $\pm$0.12 31.84 $\pm$0.09 18.63 $\pm$0.31 59.85 $\pm$0.19
ViT T5 30.33 $\pm$0.03 15.01 $\pm$0.05 66.41 $\pm$0.11 19.45 $\pm$0.34 9.22 $\pm$0.20 40.49 $\pm$0.61 31.11 $\pm$0.25 19.30 $\pm$0.45 56.24 $\pm$0.19
ImageBind 31.74 $\pm$0.21 16.45 $\pm$0.13 67.39 $\pm$0.74 24.72 $\pm$0.19 12.47 $\pm$0.09 51.32 $\pm$0.56 26.32 $\pm$0.23 16.05 $\pm$0.22 46.37 $\pm$0.66
DINOv2 T5 30.04 $\pm$0.27 14.98 $\pm$0.07 64.56 $\pm$0.56 21.77 $\pm$0.23 10.47 $\pm$0.12 45.81 $\pm$0.52 27.64 $\pm$0.95 16.21 $\pm$0.65 51.46 $\pm$1.71
MGAT CLIP 27.56 $\pm$0.30 13.55 $\pm$0.29 60.21 $\pm$0.21 21.38 $\pm$0.23 10.39 $\pm$0.22 44.60 $\pm$0.36 74.75 $\pm$1.23 64.53 $\pm$1.48 92.81 $\pm$0.64
ViT T5 30.15 $\pm$0.34 15.28 $\pm$0.34 64.84 $\pm$0.41 20.59 $\pm$0.41 9.79 $\pm$0.30 43.44 $\pm$0.76 75.26 $\pm$1.21 65.23 $\pm$1.62 92.90 $\pm$1.89
ImageBind 30.15 $\pm$0.12 15.50 $\pm$0.05 64.20 $\pm$0.43 22.13 $\pm$0.27 10.96 $\pm$0.15 45.84 $\pm$0.57 74.77 $\pm$0.49 64.95 $\pm$0.61 92.51 $\pm$0.47
DINOv2 T5 28.91 $\pm$0.09 14.47 $\pm$0.18 62.11 $\pm$0.22 21.42 $\pm$0.13 10.38 $\pm$0.13 44.11 $\pm$0.50 74.89 $\pm$1.46 64.70 $\pm$1.98 92.92 $\pm$0.41
@@
Visual encoders. For visual encoders, we select CLIP [31], ViT [5], ImageBind [10], and DINOv2 [29]. ViT and DINOv2 represent two distinct approaches to visual feature learning: ViT is explicitly trained with supervision, while DINOv2 learns robust visual features through self-supervised methods. CLIP and ImageBind are chosen to explore text-visual representation alignment. Notably, ImageBind demonstrates potential for extending multimodal graph learning across diverse modalities like audio and video by embedding them in a unified space.
Our selection of feature encoders enables us to investigate critical multimodal graph learning design, specifically:
  • The importance of aligning multiple modalities into a unified embedding space.
  • Comparative performance of supervised versus unsupervised feature encoders.
@@ 4.4 Dataloader and Evaluators
MM-Graph also provides a standardized dataloader and evaluator implementation. The dataloader is built upon PyTorch’s DataLoader class. The evaluator is based on PyTorch’s evaluator and performs standardized and reliable evaluation. The dataloader supports various sampling strategies, such as neighbor sampling and layer-wise sampling, to scale the training process to large graphs. It also enables distributed training using PyTorch’s distributed data parallel (DDP) for faster training on multiple GPUs.
By standardizing the GNNs, KGEs, feature encoders, dataloader and evaluators, MM-Graph ensures a fair and comprehensive evaluation of graph learning algorithms on multimodal graph data.
@@ 5 Empirical Analysis
We conduct a comprehensive evaluation of our proposed benchmarks on all tasks. Our experiments aim to answer the following research questions:
(RQ1) What are the performances of MM-GNNs, GNNs, and MM-KGEs on MM-Graph?
(RQ2) Do multimodal-GNNs consistently outperform conventional GNNs on MM-Graph?
(RQ3) What is the most effective approach for encoding features, and is alignment between modalities necessary?
(RQ4) How much performance gain can be achieved by using multimodal features compared to unimodal features alone?
Figure 1: (RQ2) Multimodal GNNs underperforms conventional GNNs. We compare the best performance of multimodal GNNs (MMGCN/MGAT) and conventional GNNs (SAGE, GCN, BUDDY). Conventional GNNs consistently perform better across datasets, which justifies the importance of building MM-Graph  and calls for better multimodal GNN designs.
To ensure fair and rigorous comparisons between various feature encoders and GNNs, we adopt consistent experimental settings across all experiments. We perform automatic hyperparameter tuning using Optuna [1] to optimize the performance of each model. Detailed experimental setup, hyperparameters and additional experiment results are provided in the Appendix.
@@ 5.1 Results
(RQ1) The detailed link prediction results are shown in Tab. 5. The detailed node classification results are shown in Tab. 7. The detailed knowledge graph completion results are shown in Tab. 6.
Table 6: KGC results on MM-Codex-S and MM-Codex-M. Highlighted box indicates the best performing combination for each dataset. Aligned feature embeddings outperforms unaligned features.
KGEs Dataset Image Encoder Text Encoder MRR H@1 H@3 H@10
MoSE MM-CoDEx-s CLIP CLIP 36.59 $\pm$0.16 28.10 $\pm$0.51 41.79 $\pm$1.03 51.08 $\pm$0.29
T5 ViT 39.00 $\pm$0.35 34.40 $\pm$0.47 42.73 $\pm$0.33 46.53 $\pm$0.49
ImageBind ImageBind 22.45 $\pm$1.36 5.78 $\pm$1.33 35.48 $\pm$1.30 44.49 $\pm$0.98
T5 DINOv2 40.07 $\pm$0.24 35.82 $\pm$1.85 42.77 $\pm$1.17 47.30 $\pm$0.17
MM-CoDEx-m CLIP CLIP 6.97 $\pm$0.58 3.24 $\pm$1.11 9.33 $\pm$0.33 12.15 $\pm$0.02
T5 ViT 6.09 $\pm$0.01 1.71 $\pm$0.22 8.42 $\pm$0.26 12.56 $\pm$0.28
ImageBind ImageBind 7.01 $\pm$0.15 2.26 $\pm$0.33 9.60 $\pm$0.26 13.98 $\pm$0.24
T5 DINOv2 6.40 $\pm$0.18 2.14 $\pm$0.22 9.06 $\pm$0.23 12.74 $\pm$0.14
VISTA MM-CoDEx-s CLIP CLIP 28.49 $\pm$0.39 19.12 $\pm$0.64 30.53 $\pm$0.46 49.83 $\pm$0.31
T5 ViT 29.70 $\pm$0.51 20.00 $\pm$0.59 32.11 $\pm$0.31 50.52 $\pm$0.75
ImageBind ImageBind 30.39 $\pm$0.65 20.20 $\pm$0.84 33.52 $\pm$0.30 52.68 $\pm$0.50
T5 DINOv2 26.98 $\pm$1.68 17.81 $\pm$1.16 29.53 $\pm$2.58 47.28 $\pm$2.86
MM-CoDEx-m CLIP CLIP 22.19 $\pm$1.92 15.81 $\pm$1.62 24.31 $\pm$2.10 34.73 $\pm$2.58
T5 ViT 22.10 $\pm$1.61 15.92 $\pm$1.33 24.09 $\pm$1.71 34.68 $\pm$2.26
ImageBind ImageBind 23.20 $\pm$1.45 16.97 $\pm$1.35 24.95 $\pm$1.24 35.88 $\pm$1.91
T5 DINOv2 21.38 $\pm$1.17 15.35 $\pm$1.15 23.19 $\pm$1.06 33.40 $\pm$1.49
Here are our findings:
(RQ2) Multimodal GNNs perform poorly, while GNNs and MM-KGES often outperforms on the proposed MM-Graph datasets. Our experiments, as illustrated in Figure 1 and Tab. 5, Tab. 7, and Tab. 6, reveal an unexpected trend: multimodal GNNs specifically designed for processing multimodal input data, such as MMGCN and MGAT, do not consistently outperform conventional GNN models like SAGE across various tasks and datasets. This counterintuitive result may be attributed to the architectural designs of MMGCN and MGAT, which primarily perform message passing and aggregation on each modality separately, only fusing information across modalities at the final stage. We hypothesize that this approach leads to insufficient integration or alignment of multimodal information, highlighting the need for more sophisticated model designs for multimodal graphs.
Figure 2: (RQ3) Feature alignment is important. We compare the performance of various feature encoders, find aligned features, e.g., CLIP and Imagebind result in much better performance compared with unaligned features on Amazon-Sports and Amazon-Cloth. Among them, Imagebind performs the best across backbones. This indicates the importance of using aligned features on these datasets.
In contrast, we observed that MoSE and VISTA perform notably well on the two proposed MMKG datasets, with the exception of MoSE on MM-CoDEx-m. MoSE’s approach of learning separate KGE embeddings for each modality and then ensembling them together represents a form of late fusion. While this method shows promise, its performance variability across datasets suggests that the effectiveness of late fusion strategies may be dataset-dependent.
Our new findings underscore the complexity of effectively leveraging multimodal information in graph learning tasks. They also highlight the need for further research into more advanced fusion techniques that can better capture and integrate the complementary information provided by different modalities throughout the learning process, rather than relying on late-stage integration.
(RQ3) Feature alignment is important for multimodal graph learning. Our evaluation of four feature encoders, CLIP, ImageBind, T5 + ViT, and T5 + DINOv2, reveals significant performance variations across datasets and tasks. CLIP and ImageBind, which map features from various modalities to a shared embedding space, consistently outperform T5 + ViT and T5 + DINOv2, which employ independent embedding methods for each modality without specific alignment layers.
Table 7: Node classification results on Goodreads-NC and Ele-Fashion. Conventional GNNs such as SAGE performs best across datasets. Aligned feature embeddings, e.g. CLIP and ImageBind outperforms unaligned features.
@@
MMGCN CLIP CLIP 86.10 $\pm$0.50 83.29 $\pm$0.20
T5 ViT 82.39 $\pm$0.30 81.85 $\pm$0.22
ImageBind ImageBind 86.21 $\pm$0.94 80.58 $\pm$1.08
T5 DINOv2 85.53 $\pm$0.33 82.44 $\pm$0.11
MGAT CLIP CLIP 84.66 $\pm$0.29 76.48 $\pm$0.59
T5 ViT 84.01 $\pm$0.08 75.43 $\pm$0.76
ImageBind ImageBind 86.12 $\pm$0.08 69.45 $\pm$6.25
T5 DINOv2 84.54 $\pm$0.27 74.98 $\pm$1.23
@@
Figure 3: (RQ4) Multimodal features are helpful for graph learning. Multimodal features perform better than text-only features across datasets and tasks, which justifies the necessity of introducing multimodal graph datasets.
For link prediction, aligned multimodal features consistently outperform unaligned features, as illustrated in Figure 2. ImageBind yields the best results across all backbones. However, on the Goodreads-LP dataset, T5+DINOv2 demonstrates superior performance due to its robust Optical Character Recognition (OCR) capabilities, which are crucial for extracting textual information from book covers. In node classification tasks, ImageBind and CLIP again emerge as the top-performing feature encoders. This further emphasizes the importance of feature alignment in graph-based tasks. Similarly, for knowledge graph completion, ImageBind consistently demonstrates superior performance across datasets and various KGEs, except for MM-CoDEx-s on MoSE. For MM-CoDEx-s on MoSE, we suspect that the modality split function in MoSE leads to unaligned embedding space for KGEs and thus hurts the performance.
On Goodreads-LP, the T5 + DINOv2 combination demonstrates superior performance. We believe it is because the Goodreads dataset presents a distinctive challenge because book covers primarily contain textual information, requiring robust capabilities of extracting and understanding text from book covers. This is crucial because author information is important when predicting reading preferences, and it is only displayed on book covers. Users tend to prefer books by the same author, making OCR-capable models particularly effective in this context. Our findings align with Tong et al. [40], where CLIP-based VLM is shown to be suboptimal at capturing details.
(RQ4) Multimodal models outperform unimodal models, underscoring the importance of visual information. To highlight the necessity of multimodal graph benchmarks beyond existing text-rich graphs, we conducted a study comparing the performance gains achieved by multimodal features over unimodal text or visual features alone (T5/ViT). Using an MLP model with optimal hyperparameters selected through Optuna, we focused on isolating the effects of feature types. As shown in Figure 3, integrating multimodal features consistently results in more than a 6% improvement in performance across various datasets and tasks. While visual information alone does not surpass text in performance, the combination of both modalities significantly outperforms any single modality. This finding underscores the critical role of visual information in enhancing graph representations. Note that we do not report results for knowledge graph completion (KGC) tasks, as all evaluated KGEs require multimodal inputs.
@@ 6 Conclusion
This paper presents MM-Graph, a comprehensive benchmark for multimodal graph learning. MM-Graph includes seven diverse datasets with rich textual and visual features, standardizing GNN architectures, KGEs, feature encoders, dataloaders, and evaluators to ensure fair evaluation. This setup provides insights into the performance of various GNN and encoder combinations in multimodal settings. While MM-Graph advances the integration of text and visual data, challenges remain in incorporating audio and video features due to the scarcity of open-source datasets. Although ImageBind features offer a basis for future research, further exploration is needed to effectively integrate these modalities. MM-Graph aims to drive the development of robust graph learning algorithms for real-world applications by highlighting the importance of visual information and setting the stage for future innovations in multimodal graph learning.
Acknowledgments
This material is based upon work supported by a SnapChat gift, the National Science Foundation under IIS 2212143 and CAREER Grant No. IIS 1845491. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or other funding parties.